A total of 78 seats for a concert are sold, producing total revenue of $483. If seats cost either $2.50 or $10.50, how many $10.50 seats were sold?
Answer by lwsshak3(11628)

A total of 78 seats for a concert are sold, producing total revenue of $483. If seats cost either $2.50 or $10.50, how many $10.50 seats were sold?
***
let x=number of $2.50 seats were sold
78-x=number of $10.50 seats were sold
..
2.50x+10.50(78-x)=483
2.5x+819-10.50x=483
8x=336
x=42
78-x=36
number of $10.50 seats were sold=36